# A Technical Study into 0.5B Reasoning Language Models

### v2

@@ 6 Conclusion
This research investigates the capabilities of 0.5B SRLMs in the domains of mathematical reasoning and code generation, specifically addressing the inherent challenges associated with their limited capacity. We examine various training strategies, including SFT, KD, and RL, providing a comprehensive analysis of both their potential and inherent limitations. Our empirical findings indicate that while SRLMs can be significantly enhanced through these diverse training methodologies, the effectiveness of each approach varies based on the specific task and model requirements. For instance, employing SFT integrated with Chain-of-Thought (CoT) data from KD has been demonstrated to substantially improve mathematical reasoning capabilities. In contrast, RL facilitates the transfer of valuable knowledge from larger, more complex models, enabling SRLMs to leverage existing advancements. Importantly, our analysis reveals that hybrid strategies, which combine elements of these methodologies, often yield superior outcomes compared to singular approaches. This insight underscores the necessity of tailoring training practices to the specific challenges posed by the task at hand.

### v3

@@ 6 Conclusion
This research investigates the capabilities of 0.5B SRLMs in the domains of mathematical reasoning, specifically addressing the inherent challenges associated with their limited capacity. We examine various training strategies, including SFT, KD, and RL, providing a comprehensive analysis of both their potential and inherent limitations. Our empirical findings indicate that while SRLMs can be significantly enhanced through these diverse training methodologies, the effectiveness of each approach varies based on the specific task and model requirements. For instance, employing SFT integrated with Chain-of-Thought (CoT) data from KD has been demonstrated to substantially improve mathematical reasoning capabilities. In contrast, RL facilitates the transfer of valuable knowledge from larger, more complex models, enabling SRLMs to leverage existing advancements. Importantly, our analysis reveals that hybrid strategies, which combine elements of these methodologies, often yield superior outcomes compared to singular approaches. This insight underscores the necessity of tailoring training practices to the specific challenges posed by the task at hand.
@@ Appendix A Detailed Training Strategies
Appendix B Detailed Related Works
B.1 Mathematical Reasoning
Mathematical reasoning is a complex cognitive task that requires understanding, manipulating, and solving mathematical problems [Creswell2022, Drori2022, Lewkowycz2022]. For language models, tasks involving mathematical reasoning necessitate comprehending mathematical concepts, applying logical reasoning, and performing calculations to reach correct solutions [Cobbe2021, Lample2022, Frieder2023]. This includes solving word problems, proving theorems, and carrying out algebraic manipulations [Hendrycks2021]. The model must show an understanding of mathematical principles, follow multi-step reasoning processes, and generate accurate, well-structured answers [Wang2023, Gao2022, liu2023mathematical, magister2023teaching, ahn2024large, sun2024mm, lu2023mathvista].
B.2 Small Reasoning Language Models
The field of language modeling has experienced remarkable evolution over the past decade, marked by significant advancements in both methodology and application. Early language models, characterized by their relatively small size and limited capacity, struggled to capture the complexities inherent in human language. These initial models were often constrained by limited datasets, resulting in suboptimal performance across a diverse range of linguistic tasks. The landscape began to transform with the introduction of pre-training techniques, where researchers found that training on extensive corpora of text led to considerable enhancements in models’ abilities to comprehend and generate natural language [zhao2023survey, liu2024survey, hadi2023survey]. This pivotal shift heralded the era of large language models (LLMs), which rapidly became the predominant paradigm in natural language processing.
The emergence of LLMs represented a significant milestone in AI research, epitomized by models such as GPT-3 [floridi2020gpt], which boasted billions of parameters and demonstrated unprecedented proficiencies in language understanding and generation. These models exhibited remarkable capabilities across a wide array of tasks, including text completion, translation, question answering, and summarization. Their success was largely attributed to their immense size, enabling them to discern intricate patterns and relationships within vast datasets. However, the substantial computational and resource demands associated with LLMs posed significant challenges [kaddour2023challenges]; training and deploying these models required extensive computational power, considerable energy consumption, and substantial financial investment, thereby limiting accessibility for smaller research teams and organizations with constrained resources.
In light of the limitations presented by LLMs, researchers began to explore the potential of developing smaller models that could deliver comparable performance while mitigating resource requirements. This exploration led to the advent of Small Reasoning Language Models (SRLMs), typically defined as models possessing approximately 0.5-1.0 billion parameters, such as TinyLlama [zhang2024tinyllama], Qwen2.5 series [qwen2.5] and Qwen3 series [qwen3]. SRLMs aim to strike a balance between performance and efficiency, rendering them suitable for deployment in resource-constrained environments. The impetus for SRLMs development arose from several key factors. Firstly, there was an escalating demand for models capable of operating on edge devices and within environments characterized by limited computational resources [lu2024small]. Secondly, privacy concerns associated with cloud-based LLMs prompted a search for models that could be executed locally without jeopardizing sensitive data. Finally, the environmental impact of training and deploying large models became a pressing concern, making smaller models an attractive and sustainable alternative.
SRLMs have found diverse applications across numerous domains where resource efficiency and deployment flexibility are vital. Their compact size renders them ideal for utilization in mobile applications, Internet of Things devices, and other contexts with limited computational capacity [li2025efficient]. Furthermore, their reduced computational requirements facilitate faster response times and lower latency, which are critical for real-time applications. In the realm of mathematical reasoning, SRLMs have been successfully employed to develop systems capable of solving mathematical problems, assisting in educational contexts, and supporting research endeavors in mathematics. In the area of code generation, they have been instrumental in creating tools that enhance programmers’ productivity, automate routine programming tasks, and aid in the acquisition of new programming languages.
Despite their smaller size, SRLMs have demonstrated commendable performance across various tasks. Through rigorous optimization and the adoption of advanced training techniques, researchers have been able to enhance SRLMs capabilities, allowing them to approach the performance levels of their larger counterparts. This progress has unveiled new opportunities for applying AI in scenarios where resource constraints previously posed significant barriers. However, the smaller size of SRLMs also imposes inherent limitations on their ability to learn complex patterns and relationships within data. Such constraints may hinder their performance on tasks requiring deep understanding and sophisticated reasoning, particularly in advanced mathematical problem-solving [ranaldi2024aligning] and intricate code generation [souza2025code, sqlr1]. To address these challenges, researchers have developed a range of enhancement techniques, including supervised fine-tuning, reinforcement learning [dang2025reinforcement], and knowledge distillation [chen2025sft]. These methodologies aim to maximize the potential of SRLMs while expanding their capabilities.
The objective of this research is to investigate these enhancement techniques and their applications to 0.5 billion parameter models within the specific domains of mathematical reasoning. By tackling the challenges inherent in SRLMs and capitalizing on the opportunities they present, we aspire to contribute to the advancement of efficient, specialized AI models that deliver high performance while minimizing resource requirements.
Appendix C Additional Experiment Settings
C.1 Implementation Settings
C.2 Environment